# A New Perspective on Double-S Curve Motions of Higher Order and Optimal Motion Planning

### v2

@@ 5.3 Snap-Controlled Motions
Following [6], the derivative of the jerk is called snap. We introduce $p_{\max}=w_{4}$ and $p=w_{4}=\max\limits_{t\in[0,T]}|x^{(4)}(t)|$. The equation of Theorem 5 read now
@@
allows us to plot a three-dimensional fish diagram (see Figure 8). The contour surfaces of constant $\hat{T}$ values are indicated by the colored curves circulating around the fish. The conditions
@@
define the three boundary surfaces. Combining two of these conditions yields the red boundary curves. And the combination of all the three conditions fixes the global minimum located at
@@
Figure 8: Three-dimensional plot of $\hat{T}$ as a function of $\hat{v}$, $\hat{a}$ and $\hat{\jmath}$ according to Equations (61, 65). The surfaces of constant $\hat{T}$ value are indicated by the colored curves (from orange for high $\hat{T}$ to dark purple for low $\hat{T}$), where the actual surface embraced by the curves is not shown for a better visibility of the entire figure. The three red curves frame the region of all admissible combinations of ($\hat{v},\hat{a},\hat{\jmath}$). They are obtained by combining two of the three conditions of (66), e.g., $\hat{T}_{1}=2\hat{T}_{2}$ and $\hat{T}_{2}=2\hat{T}_{3}$ fix the red solid curve. A single condition like $\hat{T}_{1}=2\hat{T}_{2}$ corresponds to a surrounding surface (not shown). The global minimum which is determined by the simultaneous consideration of all conditions of (66) is marked by the blue cross.
@@ 5.4 Systematics for Arbitrary Order
with at least 4 phases and at most 7 phases, the next order of snap-controlled motions with
@@
Also even higher order, for example the yeet $x^{(6)}(t)$ and the pow $x^{(7)}(t)$, can be considered ([18] goes to 9th order.). We will continue with an arbitrary order $N$. The dimensional transformation to obtain $\hat{x}_{0}=\hat{x}_{N}=1$ is in general given by
@@
$\hat{x}^{\star}_{n}$ $=\hat{x}_{N-1}^{{\star}{N-n}}2^{\frac{1}{2}(n-N+1)(n-N+2)}$ (76)
@@ 4 $\sqrt[4]{512}$ $\sqrt[4]{\frac{1}{32}}$ $\sqrt{\frac{1}{8}}$ $\sqrt[4]{\frac{1}{8}}$
$\hat{T}(\hat{z})=\frac{1}{\hat{z}_{1}}+\frac{\hat{z}_{1}}{\hat{z}_{2}}+\ldots+\frac{\hat{z}_{N-2}}{\hat{z}_{N-1}}+\hat{z}_{N-1}$ (84)
for $\hat{z}=(\hat{z}_{1},\ldots,\hat{z}_{N-1})\in\mathbb{R}^{N-1}$ (i.e. the time horizon) has no local minimum within the admissible region framed by the hypersurfaces encoded in $\hat{T}_{n-1}=2\hat{T}_{n}$. Thus, the solution of MTT is located at boundaries $\hat{w}_{n}$ or at surfaces $\hat{T}_{n-1}=2\hat{T}_{n}$. So $\hat{x}_{n}^{\star}=\hat{w}_{n}$ (active constraint) or there is a condition $\hat{T}_{n-1}=2\hat{T}_{n}$ for determining all $\hat{x}^{\star}_{n}$ which are not equal to a bound (i.e. $\hat{x}^{\star}_{n}<\hat{w}_{n}$). The formal algorithm to find the minimum-time trajectory reads as follows:
  1. 1.
    Transform the bounds $w_{n}$ to $\hat{w}_{n}$ according to Equation (73) within $x_{N}=w_{N}$ for $n=1,\ldots,N-1$. Set $\hat{w}_{0}=1$.
  2. 2.
    For $n=0,\ldots,N-1$ compute
    $\hat{T}_{n,\hat{w}}=\sum^{N-1}_{k=n}\frac{\hat{w}_{k}}{\hat{w}_{k+1}}.$ (85)
  3. 3.
    Collect all indices $n$ where $\hat{T}_{n-1,\hat{w}}<2\hat{T}_{n,\hat{w}}$, i.e. find the sets
    $B$ $=\left\{n\penalty 10000\ |\penalty 10000\ \hat{T}_{n-1,\hat{w}}<2\hat{T}_{n,\hat{w}}\wedge 0\leq n\leq N-1\right\}$ (86)
    $\bar{B}$ $\B\displaystyle=\left\{0,\ldots,N-1\right\}\backslash$ B (87)
  4. 4.
    For all $n\in\bar{B}$ set $\hat{x}^{\star}_{n}=\hat{w}_{n}$.
  5. 5.
    Solve the equation system
[…]
    for $\hat{z}\in\mathbb{R}^{n}$ with $n\in B$ and $\hat{T}_{n}(\hat{z})=\sum_{k=n}^{N-1}\frac{\hat{z}_{k}}{\hat{z}_{k-1}}$ where $\hat{z}_{0}=\hat{z}_{N}=1$ and $\hat{z}_{n}=\hat{x}^{\star}_{n}=\hat{w}_{n}$ for $n\in\bar{B}$.
This algorithm allows a direct computation of the reached values of velocity, acceleration, $\ldots$, without complicated case distinction. Since the $\hat{T}_{n}$ values follow directly the trajectory can be constructed directly as well.
@@ 6 Summary and Outlook
for jerk-controlled motions, is valid for any number of phases the motion can have. Furthermore, arbitrarily high orders can be considered. This allows for a reinterpretation and solution of well-known standard trajectory optimization problems (such as motions with minimum time or motions with minimum velocity). In particular, the condition $T_{n}=2T_{n+1}$ enables the determination of the active constraints for the respective case. It can also be estimated what effect considering higher orders has on the overall time horizon. Therefore, it seems surprising that – although the equation is not exceptionally complex and covers a large number of cases – it has not previously been presented in this form in the literature.
@@
Clearly, many extensions exist. First, the zero boundary condition can be relaxed. Furthermore, it would be possible to consider other classes of trajectories (for example trajectories with circular blends of trigonometric trajectories). The expected results would then likely be considerably more unwieldy than those obtained here.

### v3

@@ 5.3 Snap-Controlled Motions
Following the terminology of [6], the derivative of the jerk is called snap. We introduce $p_{\max}=w_{4}$ and $p=w_{4}=\max\limits_{t\in[0,T]}|x^{(4)}(t)|$. Equation (32) of Theorem 5 reads now
@@
allows us to plot a three-dimensional diving fish diagram (see Figure 9). The contour surfaces of constant $\hat{T}$ values are indicated by the colored curves circulating around the fish. The conditions
@@
define the three boundary surfaces. Combining two of these conditions yields the red boundary curves. And the combination of all the three conditions fixes the global minimum (see Lemma 9) located at
@@
Figure 9: Three-dimensional plot of $\hat{T}$ as a function of $\hat{v}$, $\hat{a}$ and $\hat{\jmath}$ according to Equations (62, 66). The surfaces of constant $\hat{T}$ value are indicated by the colored curves (from orange for high $\hat{T}$ to deep purple for low $\hat{T}$), where the actual surface embraced by the curves is not shown for a better visibility of the entire figure. The three red curves frame the region of all admissible combinations of ($\hat{v},\hat{a},\hat{\jmath}$). They are obtained by combining two of the three conditions of (67), e.g., combining $\hat{T}_{1}=2\hat{T}_{2}$ and $\hat{T}_{2}=2\hat{T}_{3}$ fixes the red solid curve. A single condition like $\hat{T}_{1}=2\hat{T}_{2}$ corresponds to a surrounding surface (not shown). The global minimum which is determined by the simultaneous consideration of all conditions of (67) is marked by the blue cross.
@@ 5.4 Systematics for Arbitrary Order
with at least 4 phases and at most 7 phases, the next order of snap-controlled motions obeys
@@
Also even higher orders, for example the yeet $x^{(6)}(t)$ and the pow $x^{(7)}(t)$, can be considered ([24] goes to 9th order). We will continue with an arbitrary order $N$. The dimensional transformation to obtain $\hat{x}_{0}=\hat{x}_{N}=1$ is in general given by
@@
$\hat{x}^{\star}_{n}$ $=\hat{x}_{N-1}^{{\star}{N-n}}2^{\frac{1}{2}(n-N+1)(n-N+2)}=2^{\frac{n}{2}(\frac{2}{N}-N+n)}$ (77)
@@ 4 $\sqrt[4]{512}$ $\sqrt[4]{\frac{1}{32}}$ $\sqrt{\frac{1}{8}}$ $\sqrt[4]{\frac{1}{8}}$
5.5 Computation of Minimum-Time Trajectories
Let us eventually consider MTT algorithmically. Although not crucial, we will mainly work with dimensionless quantities. It is convenient to define the earliest time point of achievement $T_{n}$ (see Definition 4) as a function on $\mathbb{R}^{N+1}$: For $z=(z_{0},\ldots,z_{N})\in\mathbb{R}^{N+1}$ set
$T_{n}(z)=\sum_{k=n}^{N-1}\frac{z_{k}}{z_{k+1}}.$ (85)
The main idea [^11] Note that minimizing the time horizon is a convex optimization problem since the objective function $T_{0}$ is convex and the convex combination of two admissible trajectories is admissible as well (cf. [26] for instance). is based on the observation that the function $T_{0}(z)$ attains no inner minimum [^12] Regarding the entire plane $\mathbb{R}^{N+1}$, it is a straight forward calculation to show that $z_{0}=\ldots=z_{N}=1$ is the global minimum. within the admissible region framed by the hypersurfaces encoded in $T_{n-1}(z)=2T_{n}(z)$. Therefore, either a bound $w_{n}$ is active, or the equation $T_{n-1}(z)=2T_{n}(z)$ holds.
Let a distance $s$ as well as the bounds $w_{1},\ldots,w_{N}$ be given. Then the steps of the algorithm for finding the minimum-time motion are as follows:
  1. 1.
    Transform $w_{n}$ to $\hat{w}_{n}$ according to Equation (74), and set $\hat{u}=(\hat{u}_{0},\ldots,\hat{u}_{N})$ with
    $\hat{u}_{n}=\begin{cases}1&n=0,N\\ \min\left\{\hat{w}_{n},\hat{x}_{n}^{\star}\right\}&0<n<N\end{cases}$ (86)
  2. 2.
    Collect all indices $1\leq n\leq N-1$ with $T_{n-1}(\hat{u}_{n})<2T_{n}(\hat{u}_{n})$ (see Equation (85)), i.e. find the set
    $B=\left\{n\penalty 10000\ |\penalty 10000\ T_{n-1}(\hat{u}_{n})<2T_{n}(\hat{u}_{n})\wedge 1\leq n\leq N-1\right\}.$ (87)
  3. 3.
    Solve [^13] E.g., with Newton’s method. the system of equations
    $T_{n-1}(\hat{z})=2T_{n}(\hat{z})\quad(n\in B)$ (88)
    for $\hat{z}=(\hat{z}_{0},\ldots,\hat{z}_{N}),\hat{z}_{0}=\hat{z}_{N}=1$ including $\hat{z}_{n}=\hat{u}_{n}$ for $n\notin B$.
  4. 4.
    Perform a back transformation from $\hat{z}$ to gain the solution $x_{\mathrm{MTT}}$ and the time points of earliest achievement as $T_{n}(x_{\mathrm{MTT}})$.
Let us illustrate the algorithm by the following example of 5th order. We seek for the minimum time motion accomplishing $s=20$ m subject to constraints
$w_{1}$ $=v_{\max}=7\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}}$
$w_{2}$ $=a_{\max}=2\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{2}}$
$w_{3}$ $=j_{\max}=0.5\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{3}}$
$w_{4}$ $=p_{\max}=6\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{4}}$
$w_{5}$ $=c_{\max}=10\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{5}}.$
Table 2 summarizes the steps to determine the minimum-time motion. The ratio $\frac{T_{2}(\hat{u})}{T_{3}(\hat{u})}>2$ fixes $\hat{z}_{3}=\hat{w}_{3}=0.0379$ (and $j_{\mathrm{MTT}}=j_{\max}=0.5\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{3}}$). For the remaining quantities $\hat{z}_{n}$ we have to solve
$\frac{1}{\hat{z}_{1}}$ $=\frac{\hat{z}_{1}}{\hat{z}_{2}}+\frac{\hat{z}_{2}}{\hat{w}_{3}}+\frac{\hat{w}_{3}}{\hat{z}_{4}}+\hat{z}_{4}$ (89)
$\frac{\hat{z}_{1}}{\hat{z}_{2}}$ $=\frac{\hat{z}_{2}}{\hat{w}_{3}}+\frac{\hat{w}_{3}}{\hat{z}_{4}}+\hat{z}_{4}$ (90)
$\frac{\hat{w}_{3}}{\hat{z}_{4}}$ $=\hat{z}_{4}$ (91)
$s$ $v_{\max}$ $a_{\max}$ $j_{\max}$ $p_{\max}$ $c_{\max}$
$w_{1}$ $w_{2}$ $w_{3}$ $w_{4}$ $w_{5}$
20 m $7\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}}$ $2\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s^{2}}}$ $0.5\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s^{3}}}$ $6\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s^{4}}}$ $10\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s^{5}}}$
step 1 $\hat{w}_{1}$ $\hat{w}_{2}$ $\hat{w}_{3}$ $\hat{w}_{4}$
0.4020 0.1320 0.0379 0.5223
$\hat{u}_{0}$ $\hat{u}_{1}$ $\hat{u}_{2}$ $\hat{u}_{3}$ $\hat{u}_{4}$ $\hat{u}_{5}$
1 0.2873 0.1320 0.0379 0.4353 1
@@ 5.5 Computation of Minimum-Time Trajectories
9.6631 6.1809 4.0045 0.5223 0.4353
ratio ratio ratio ratio
1.56 1.54 7.67 1.20
step 3 $\hat{z}_{0}$ $\hat{z}_{1}$ $\hat{z}_{2}$ $\hat{z}_{3}$ $\hat{z}_{4}$ $\hat{z}_{5}$
1 0.1999 0.0799 0.0379 0.1947 1
step 4 $x_{\mathrm{MTT},0}$ $x_{\mathrm{MTT},1}$ $x_{\mathrm{MTT},2}$ $x_{\mathrm{MTT},3}$ $x_{\mathrm{MTT},4}$ $x_{\mathrm{MTT},5}$
20 m 3.48 $\frac{\mathrm{m}}{\mathrm{s}}$ 1.21 $\frac{\mathrm{m}}{\mathrm{s^{2}}}$ 0.5 $\frac{\mathrm{m}}{\mathrm{s^{3}}}$ 2.24 $\frac{\mathrm{m}}{\mathrm{s^{4}}}$ 10 $\frac{\mathrm{m}}{\mathrm{s^{5}}}$
$T_{0}$ $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$
11.49 s 5.74 s 2.87 s 0.45 s 0.22 s
Table 2: Summary of exemplary time minimization for a distance $s=20$ m and the above constraints. The minimal time is $T=8.4481$ s.
where we already inserted $\hat{w}_{3}$ as $\hat{z}_{3}$. Starting from (91), the solution is not difficult to find (see Table 2). We have $\hat{T}_{0}=10.0$ corresponding to $T=T_{0}=11.49\penalty 10000\ \mathrm{s}$ as minimal time. Figure 10 shows all orders of the optimal trajectory. Note that there are two interesting cases when varying the bound of the highest order $w_{5}=c_{\max}$ (see Figure 11):
  • If $c_{\max}=72\penalty 10000\ \frac{\mathrm{m}}{\mathrm{s}^{5}}$, then both $j_{\max}$ and $p_{\max}$ are always reached. Independent of $c_{\max}$, $a_{\max}$ and $v_{\max}$ can never be reached since $s$ is too short.
  • If $c_{\max}\leq 0.683$, then no bound is reached anymore and the solution of MTT is equivalent to the global minimum $x_{1}^{\star},\ldots,x^{\star}_{N-1}$ (see Lemma 9). Graphically speaking, the small value of $c_{\max}$ catapults the transformed boundaries $\hat{w}_{n}$ to high values lying for beyond the tail of the fish.
In Appendix A we provide an implementation of the algorithm in Python and one implementation in C++. Solving the above example with the Python implementation [^14] Note that we did not place much importance to the best possible high-end implementation. For instance, the Python implementation could be enhanced through a modification of Newton’s method by incorporating analytic derivatives instead of using numerical derivatives. took about 1 ms on an Intel® Core i5-8265U.
This algorithm allows a direct computation of the reached values of velocity, acceleration, etc. Since the values $T_{n}$ follow directly the trajectory can be constructed directly as well.
Figure 10: Dimensionless profiles of the minimum-time trajectory for $s=20$ m (see Table 2), where $\hat{x}_{0}$ show the position as a function of the time (higher orders accordingly). The dimensionless time horizon of the motion is $\hat{T}_{0}=10.00$ (equivalent to 11.49 s). The maximal velocity is initially achieved at $\hat{T}_{1}=5.00$ (5.74 s) – and higher orders analogously.
Figure 11: Time horizon of the minimum-time motion as a function of $c_{\max}$. The other parameters are of the example displayed in Figure 10.
Finally, with the insights gained from the above algorithm, the MDT optimization can be treated similarly. Let us sketch the procedure. Recall that a distance $s$ and a time horizon $T_{\mathrm{MDT}}$ are given. The objective function is $x_{M}=\max\limits_{t\in[0,T]}\left|x^{(M)}(t)\right|\rightarrow\min$ for $1\leq M<N$. As a first step, $T_{M}$ has to be minimized. Observe that the smaller $T_{M}$ the longer $x_{M}$ can be held and consequently the smaller becomes the objective function. The minimization of $T_{M}$ can be done by applying the above algorithm to the reduced problem of the orders $n=N\ldots M$. This yields a part of the solution, namely $x_{\mathrm{MDT},N}$, …, $x_{\mathrm{MDT},M}$. It remains to find $x_{\mathrm{MDT},M-1}$, …, $x_{\mathrm{MDT},1}$. To this end, we have to proceed analogously as above but with two differences: first, the part of the solution for $n\geq M$ is already determined by the $T_{M}$ minimization. Second, $T_{0}$ is fixed as $T_{\mathrm{MDT}}$.
@@ 6 Summary and Outlook
for jerk-controlled motions, is valid for any number of phases the motion can have. Furthermore, arbitrarily high orders can be considered. For example,
$T=\frac{s}{v}+\frac{v}{a}+\frac{a}{j}+\frac{j}{p}+\ldots=\sum_{n=0}^{N-1}\frac{x_{n}}{x_{n+1}}.$ (93)
This allows for a reinterpretation and solution of well-known standard trajectory optimization problems (such as motions with minimum time or motions with minimum velocity). In particular, the condition
$T_{n}=2T_{n+1}$ (94)
enables the determination of the active constraints for the respective case. Additionally, the effect of incorporating higher orders on the time horizon can be considered as well. Therefore, it seems surprising that – although the equation is not exceptionally complex and covers a large number of cases – it has not previously been presented in this form in the literature.
@@
Clearly, many extensions exist. First, the zero boundary condition can be relaxed. Furthermore, it would be possible to consider other classes of trajectories (for example trajectories with circular blends of trigonometric trajectories). The expected results would then likely be considerably more cumbersome than those obtained here.
@@ Acknowledgement
Appendix A Implementations of the Optimization Algorithm
Both implementations below take the transformed boundaries as input and return the solution as dimensionless quantities.
A.1 Python Implementation
import numpy as np
from scipy.optimize import root
# Step 1: computation of vector u
def compute_u(N, w):
w = np.asarray(w, dtype=float)
u = np.zeros(N + 1, dtype=float)
u[0] = 1.0
u[N] = 1.0
for n in range(1, N):
exponent = 0.5 * n * (2.0 / N - N + n)
bound = 2 ** exponent
u[n] = min(w[n - 1], bound)
return u
[…]
def T(n, x):
N = len(x) - 1
return np.sum(x[n:N] / x[n+1:N+1])
# Step 2: determination of the set B
def compute_B(z):
N = len(z) - 1
B = []
for n in range(1, N):
if T(n - 1, z) < 2 * T(n, z):
B.append(n)
return np.array(B, dtype=int)
# Step 3: setting up the equation system
def residuals(x, z, B):
N = len(x) - 1
eqs = []
B = set(B)
# x_0 = x_N = 1
eqs.append(x[0] - 1)
eqs.append(x[N] - 1)
# equations for n=1, …, N-1
for n in range(1, N):
if n in B:
eqs.append(T(n - 1, x) - 2*T(n, x)) # n in B
else:
eqs.append(x[n] - z[n]) # n not in B
return np.array(eqs)
# main algorithm
def solve_system(N, w, x0=None):
u = compute_u(N, w)
B = compute_B(u)
if x0 is None:
x0 = u.copy()
sol = root(residuals, x0, args=(u, B), method=’hybr’)
if not sol.success:
raise RuntimeError("no solution", sol.message)
x = sol.x
if np.any(x <= 0):
raise RuntimeError("solution not positive")
t = np.array([T(n, x) for n in range(0, N)])
return x, t, B, u
# example
if __name__ == "__main__":
N = 5
w = np.array([0.402044,0.131951,0.037893,0.522330])
x, t, B, u = solve_system(N, w)
print("u =", u)
print("B =", B)
print("x =", x)
print("t =", t)
A.2 C++ Implementation
#include <iostream>
#include <vector>
#include <cmath>
#include <stdexcept>
#include <iomanip>
using namespace std;
[…]
double T(int n, const vector<double>& x) {
int N = x.size() - 1;
double s = 0.0;
for (int k = n; k <= N - 1; k++) {
s += x[k] / x[k + 1];
return s;
// Step 1: computation of vector u
vector<double> compute_u(int N, const vector<double>& w) {
vector<double> u(N + 1);
u[0] = 1.0;
u[N] = 1.0;
for (int n = 1; n <= N - 1; n++) {
double exponent = 0.5 * n * (2.0 / N - N + n);
double bound = pow(2.0, exponent);
u[n] = min(w[n - 1], bound);
return u;
// Step 2: determination of the set B
vector<int> compute_B(const vector<double>& z) {
int N = z.size() - 1;
vector<int> B;
for (int n = 1; n <= N - 1; n++) {
if (T(n - 1, z) < 2.0 * T(n, z))
B.push_back(n);
return B;
// Newton method: computation of the residuals
vector<double> residuals(const vector<double>& x,
const vector<double>& z,
const vector<int>& Bset) {
int N = x.size() - 1;
vector<double> r(N + 1, 0.0);
// B as boolean index
vector<bool> inB(N + 1, false);
for (int b : Bset) inB[b] = true;
// x_0 = x_N = 1
r[0] = x[0] - 1.0;
r[N] = x[N] - 1.0;
// equations
for (int n = 1; n <= N - 1; n++) {
if (inB[n]) { // n in B
r[n] = T(n - 1, x) - 2.0 * T(n, x);
} else { // n not in B
r[n] = x[n] - z[n];
return r;
// Newton method for solving the equation system
vector<double> solve_nonlin(const vector<double>& z, const vector<int>& B) {
int N = z.size() - 1;
vector<double> x = z; // initial value: u
const double eps = 1e-10;
const double h = 1e-6;
const int maxIter = 200;
for (int it = 0; it < maxIter; it++) {
vector<double> F = residuals(x, z, B);
// checking for convergence
double norm = 0.0;
for (double v : F) norm += v * v;
norm = sqrt(norm);
if (norm < eps) {
return x;
// numerical Jacobian
vector<vector<double>> J(N + 1, vector<double>(N + 1, 0.0));
for (int j = 0; j <= N; j++) {
vector<double> xh = x;
xh[j] += h;
vector<double> Fh = residuals(xh, z, B);
for (int i = 0; i <= N; i++) {
J[i][j] = (Fh[i] - F[i]) / h;
// solving linear system J * dx = -F with Gauss
vector<double> dx = F;
for (double &v : dx) v = -v;
// Gauss elimination:
for (int i = 0; i <= N; i++) {
double piv = J[i][i];
if (fabs(piv) < 1e-14)
throw runtime_error("Newton: singular matrix.");
// normalizing
for (int j = i; j <= N; j++) J[i][j] /= piv;
dx[i] /= piv;
// elimination
for (int k = 0; k <= N; k++) {
if (k == i) continue;
double f = J[k][i];
for (int j = i; j <= N; j++) {
J[k][j] -= f * J[i][j];
dx[k] -= f * dx[i];
// next iteration point
for (int i = 0; i <= N; i++) {
x[i] += dx[i];
throw runtime_error("Newton method did not converge.");
// main algorithm and example
int main() {
int N = 5;
vector<double> w = {0.402044,0.131951,0.037893,0.522330};
vector<double> u = compute_u(N, w);
vector<int> B = compute_B(z);
vector<double> x = solve_nonlin(u, B);
// t vektor: t_n = T_n(x)
vector<double> t(N);
for (int n = 0; n < N; n++) {
t[n] = T(n, x);
// output
cout << "u = ";
for (double v : u) cout << v << " ";
[…]
for (int b : B) cout << b << " ";
[…]
for (double v : x) cout << v << " ";
[…]
for (double v : t) cout << v << " ";
cout << endl;
return 0;